# VocSegMRI: Multimodal Learning for Precise Vocal Tract Segmentation in Real-time MRI

### v2

@@ 3 Experiments and results
Table 1: Segmentation performance of different models and input modalities. ASSD/HD are given in millimeters ($m$).
[…]
Image-Only Baseline
U-Net 0.81 $\pm$ 0.06 0.89 $\pm$ 0.03 3.64 $\pm$ 0.77 8.08 $\pm$ 1.21
Swin UNETR 0.82 $\pm$ 0.05 0.89 $\pm$ 0.02 3.21 $\pm$ 0.58 7.54 $\pm$ 1.17
SAM-Med2D 0.83 $\pm$ 0.05 0.91 $\pm$ 0.02 3.23 $\pm$ 0.61 7.67 $\pm$ 1.12
ResNet-50 0.84 $\pm$ 0.04 0.91 $\pm$ 0.02 3.04 $\pm$ 0.57 7.22 $\pm$ 1.08
nnU-Net 0.86 $\pm$ 0.03 0.93 $\pm$ 0.02 1.74 $\pm$ 0.43 5.37 $\pm$ 1.08
ViT-base 0.86 $\pm$ 0.02 0.94 $\pm$ 0.01 2.11 $\pm$ 0.58 4.73 $\pm$ 0.91
Concat Fusion
V 0.86 $\pm$ 0.02 0.94 $\pm$ 0.01 2.11 $\pm$ 0.58 4.73 $\pm$ 0.91
VA 0.87 $\pm$ 0.03 0.91 $\pm$ 0.03 2.40 $\pm$ 0.49 5.81 $\pm$ 0.98
VP 0.86 $\pm$ 0.04 0.89 $\pm$ 0.02 3.21 $\pm$ 0.58 7.54 $\pm$ 1.17
VAP 0.89 $\pm$ 0.02 0.94 $\pm$ 0.01 2.19 $\pm$ 0.41 5.00 $\pm$ 0.97
Ablation Study
Cross-Att 0.90 $\pm$ 0.02 0.95 $\pm$ 0.01 1.83 $\pm$ 0.37 4.91 $\pm$ 1.00
Contrastive 0.89 $\pm$ 0.02 0.93 $\pm$ 0.01 2.03 $\pm$ 0.44 4.35 $\pm$ 0.87
VocSegMRI 0.91 $\pm$ 0.01 0.95 $\pm$ 0.01 1.52 $\pm$ 0.31 4.26 $\pm$ 0.88
V: video. A: audio. P: phonological. Cross-Att: cross-attention.
We designed three experimental groups, with results presented in the upper, middle, and lower sections of Tab. 1. Evaluation metrics include Intersection over Union (IoU), Dice coefficient, Average Symmetric Surface Distance (ASSD) and HD95. All metrics are reported as mean $\pm$ standard deviation (std). In the Image-Only Baseline, we compared the performance of several state-of-the-art (SOTA) segmentation models using only frame-wise rtMRI images as input [22, 23, 24, 25, 26, 27]. Among them, ViT-base and nnU-Net achieved the highest IoU of 0.86, followed by ResNet (0.84) and SAM-Med2D (0.83). In the Concat Fusion setting, we implemented multimodal fusion by concatenating the outputs of the respective encoders for each modality. We evaluated segmentation performance using different combinations of input modalities. Adding either audio or phonological class features led to improvements in IoU over the ViT (0.86), with varying magnitudes. The best performance was achieved using all three modalities (IoU==0.89). Finally, in the Ablation Study configuration, we evaluated the contribution of each component individually. The Cross configuration incorporates only cross-attention, while Contrast introduces only contrastive learning. Our proposed model, VocSegMRI, combines both mechanisms, achieving the best overall performance. yielding the highest Dice score (0.95) and the lowest ASSD (1.52) and HD95 (4.26).
@@
Fig. 3: Qualitative comparison of ground-truth (GT) and predicted segmentations for one rtMRI frame
@@
We selected four representative models for comparison. Fig. 3 presents qualitative segmentation results for the Tongue and Lower Lip, highlighting false positive (FP) and false negative (FN) regions, as well as reporting precision/recall values, while nnU-Net performed the worst among all models. Specifically, for the Tongue, nnU-Net achieved a precision of 0.91 and recall of 0.80, while for the Lower Lip, precision was 0.42 and recall 0.97. Incorporating all three modalities as input led to improved model performance. However, VocSegMRI consistently achieved the best trade-off across both structures, with an overall precision of 0.85 and recall of 0.98.
@@ 4 Discussion and Conclusions
Class-level analysis (Fig. 2) indicates that segmentation is more accurate for larger structures such as the Tongue and Velum, whereas smaller articulators (Upper/Lower Lips) remain challenging due to their very low pixel representation in the images and weaker visual cues. For instance, in the example shown in Fig. 3, the Lower Lip achieves only 0.14 precision with nnU-Net and 0.42 precision with the ViT baseline, despite high recall values (0.38 and 0.97, respectively), reflecting numerous FP. In contrast, the Tongue exhibits higher precision across models (0.75–0.91) with high recall. Qualitative comparisons further illustrate that unimodal models suffer from pronounced FP and FN, particularly along anatomical boundaries. While simple concatenation partially mitigates these errors, cross-attention combined with contrastive learning enables more precise and balanced segmentation, improving Lower Lip precision to 0.68 and overall precision-recall trade-off to 0.85/0.98, thereby substantially reducing both false positives and false negatives across all structures.

### v3

@@ 3 Experiments and results
We designed three experimental groups, with results presented in the upper, middle, and lower sections of Tab. Evaluation metrics include Intersection over Union (IoU), Dice coefficient, Average Symmetric Surface Distance (ASSD) and HD95. All metrics are reported as mean $\pm$ standard deviation (std). In the Image-Only Baseline, we compared the performance of several state-of-the-art (SOTA) segmentation models using only frame-wise rtMRI images as input [22, 23, 24, 25, 26, 27]. Among them, ViT-base and nnU-Net achieved the highest IoU of 0.86, followed by ResNet (0.84) and SAM-Med2D (0.83). In the Concat Fusion setting, we implemented multimodal fusion by concatenating the outputs of the respective encoders for each modality. We evaluated segmentation performance using different combinations of input modalities. Adding either audio or phonological class features led to improvements in IoU over the ViT (0.86), with varying magnitudes. The best performance was achieved using all three modalities (IoU==0.89). Finally, in the Ablation Study configuration, we evaluated the contribution of each component individually. The Cross configuration incorporates only cross-attention, while Contrast introduces only contrastive learning. Our proposed model, VocSegMRI, combines both mechanisms, achieving the best overall performance. yielding the highest Dice score (0.95) and the lowest ASSD (1.52) and HD95 (4.26).
@@
We selected four representative models for comparison. Fig presents qualitative segmentation results for the Tongue and Lower Lip, highlighting false positive (FP) and false negative (FN) regions, as well as reporting precision/recall values, while nnU-Net performed the worst among all models. Specifically, for the Tongue, nnU-Net achieved a precision of 0.91 and recall of 0.80, while for the Lower Lip, precision was 0.42 and recall 0.97. Incorporating all three modalities as input led to improved model performance. However, VocSegMRI consistently achieved the best trade-off across both structures, with an overall precision of 0.85 and recall of 0.98.
@@ 4 Discussion and Conclusions
Class-level analysis (Fig. 2) indicates that segmentation is more accurate for larger structures such as the Tongue and Velum, whereas smaller articulators (Upper/Lower Lips) remain challenging due to their very low pixel representation in the images and weaker visual cues. For instance, in the example shown in Fig, the Lower Lip achieves only 0.14 precision with nnU-Net and 0.42 precision with the ViT baseline, despite high recall values (0.38 and 0.97, respectively), reflecting numerous FP. In contrast, the Tongue exhibits higher precision across models (0.75–0.91) with high recall. Qualitative comparisons further illustrate that unimodal models suffer from pronounced FP and FN, particularly along anatomical boundaries. While simple concatenation partially mitigates these errors, cross-attention combined with contrastive learning enables more precise and balanced segmentation, improving Lower Lip precision to 0.68 and overall precision-recall trade-off to 0.85/0.98, thereby substantially reducing both false positives and false negatives across all structures.